Need asap please!Considering two firms, Alpha and Beta, the sales, profits, and assets of Alpha are $20 million, $4 million, and $40 million r
Answer #1

Requirement 1:

Return on Assets = Profit/Assets

Return on Assets for Alpha = 4/40 = 10%

Return on Assets for Beta = 2/20 = 10%

Eventhough the firms are operating at different level of assets & profits, they both have same Return on Assets %.

Requirement 2:

Return on Assets under Dupont System = Profit margin * Asset Turnover

= Profit/Sales * Sales/Assets

ROA under Dupont system for Alpha = 4/20 * 20/40 = 10%

ROA under Dupont system for Beta = 2/8 * 8/20 = 10%

By Applying Dupont we can obtain the info about profit margin & assets Turnover which can be used for comparison to other equal firms.